2. Christopher construction, Inc. has current assets of $20,000 and current liabilities of $10,000. If we assume that each transaction is independent, what is the effect of each of the following transaction on Christopher Construction Inc.’s current ratio? Compute the current ratio for each case. a) $2000 of accounts payable are paid off with cash. b) Inventories of $5000 are purchased on credit. c) Additional common stock is sold for $4000 cash. d) A long-term debt of $2,000 is obtained, and the proceeds are used to buy a new machine. e) Account receivable of $1,000 are collected in cash. 2. Christopher construction, Inc. has current assets of $20,000 and current liabilities of $10,000. If we assume that each transaction is independent, what is the effect of each of the following transaction on Christopher Construction Inc.’s current ratio? Compute the current ratio for each case. a) $2000 of accounts payable are paid off with cash. b) Inventories of $5000 are purchased on credit. c) Additional common stock is sold for $4000 cash. d) A long-term debt of $2,000 is obtained, and the proceeds are used to buy a new machine. e) Account receivable of $1,000 are collected in cash. 2. Christopher construction, Inc. has current assets of $20,000 and current liabilities of $10,000. If we assume that each transaction is independent, what is the effect of each of the following transaction on Christopher Construction Inc.’s current ratio? Compute the current ratio for each case. a) $2000 of accounts payable are paid off with cash. b) Inventories of $5000 are purchased on credit. c) Additional common stock is sold for $4000 cash. d) A long-term debt of $2,000 is obtained, and the proceeds are used to buy a new machine. e) Account receivable of $1,000 are collected in cash.
Answer #1

Current Ratio = Current Assets/Current Liabilities

= 20,000/10,000

= 2

a)$2000 of accounts payable are paid off with cash: Current liabilities will be reduced by $2,000 and current assets will also reduce by $2,000

Current Ratio = 18,000/8,000

= 2.25

b) Inventories of $5000 are purchased on credit: Inventories will increase and accounts payables will increase

Current Ratio = 25,000/15,000

= 1.67

c) Additional common stock is sold for $4000 cash: Cash increases, it means current assets will increase, no change in current liabilities

Current Ratio = 24,000/10,000

= 2.4

d) A long-term debt of $2,000 is obtained, and the proceeds are used to buy a new machine: No change in current assets and current liabilities

Current ratio = 20,000/10,000

=2

e) Account receivable of $1,000 are collected in cash: Changes in assets, reduction in accounts receivables and increase in cash, no net change in current assets

Current Ratio = 20,000/10,000

= 2